# Probabilistic Robustness Analysis in High Dimensional Space: Application to Semantic Segmentation Network

### v1

@@ Abstract
Semantic segmentation networks (SSNs) play a critical role in domains such as medical imaging, autonomous driving, and environmental monitoring, where safety hinges on reliable model behavior under uncertainty. Yet, existing probabilistic verification approaches struggle to scale with the complexity and dimensionality of modern segmentation tasks, often yielding guarantees that are too conservative to be practical. We introduce a probabilistic verification framework that is both architecture-agnostic and scalable to high-dimensional outputs. Our approach combines sampling-based reachability analysis with conformal inference (CI) to deliver provable guarantees while avoiding the excessive conservatism of prior methods. To counteract CI’s limitations in high-dimensional settings, we propose novel strategies that reduce conservatism without compromising rigor. Empirical evaluation on large-scale segmentation models across CamVid, OCTA-500, Lung Segmentation, and Cityscapes demonstrates that our method provides reliable safety guarantees while substantially tightening bounds compared to SOTA. We also provide a toolbox implementing this technique, available on github. [^1] https://github.com/Navidhashemicodes/SSN_Reach_CLP_Surrogate
@@ 1 Introduction
Deep Neural Networks have achieved outstanding success across diverse domains, including image classification, speech recognition, medical diagnosis, and autonomous systems. However, despite their effectiveness, DNNs remain highly sensitive to input perturbations and adversarial examples—carefully crafted modifications that can cause significant changes in model predictions. This vulnerability poses serious risks, particularly in safety-critical applications such as autonomous driving, and medical imaging, where even a single misprediction may lead to catastrophic consequences.
Research on neural network robustness has generally followed two main directions. Falsification-based approaches aim to uncover adversarial examples that expose a model’s vulnerabilities (Madry et al., 2017; Goodfellow et al., 2014; Kurakin et al., 2016). While effective for identifying potential failure cases, these methods provide no formal guarantees of safety. In contrast, verification-based approaches seek to certify robustness by proving that a network remains reliable under specified input perturbations (Zhou et al., 2024; Lemesle et al., 2024; Wu et al., 2024). Although these techniques offer provable guarantees, they are typically hindered by severe scalability challenges.
Reachability analysis—the task of computing the set of all possible outputs of a system given a bounded input set—has become a central tool for verifying robustness. In the context of neural networks, it captures how input perturbations propagate through the layers and affect the final prediction. However, reachability analysis quickly becomes computationally infeasible for large-scale models or high-dimensional outputs. Semantic segmentation networks (SSNs), which assign a class label to every pixel, present a particularly difficult case due to their extremely high-dimensional output space and complex architectures, often involving components such as transposed convolutions, max-pooling, and dilated convolutions. For such models, traditional reachability-based methods either fail to scale or yield overly conservative guarantees.
To overcome these challenges, probabilistic verification has emerged as a promising direction (Fischer et al., 2021; Anani et al., 2024; Hao et al., 2022; Marzari et al., 2024). Instead of requiring guarantees to hold for every possible perturbation, this approach establishes them with provable probability under a prescribed distribution of input uncertainty. Such methods tend to scale more effectively and yield informative statistical assurances, particularly when exact reachability bounds or Lipschitz estimates are impractical or overly pessimistic. Yet, current probabilistic techniques remain limited: they often cannot capture complex output specifications and struggle to handle the scale of modern semantic segmentation networks. In this paper, we introduce a scalable, architecture-independent probabilistic verification framework with reduced conservatism, specifically designed for semantic segmentation models. Our approach makes no restrictive assumptions on network layers or structure, and is naturally applicable to high-dimensional outputs.
@@
[…]
Deterministic Verification of Neural Networks. Deterministic verification of neural networks aims to provide formal guarantees that the network’s outputs meet a given specification for all inputs within a prescribed set. This problem has been addressed using several well-established techniques, such as Satisfiability Modulo Theories (SMT) solvers (Duong et al., 2023; Katz et al., 2017; Wu et al., 2024), Mixed Integer Linear Programming (MILP) formulations (Anderson et al., 2020; Cheng et al., 2017; Tjeng et al., 2017), and different types of reachability analysis (Bak et al., 2020; Tran et al., 2020). Many of these methods can be viewed through the general framework of branch-and-bound algorithms (Bunel et al., 2020), which iteratively partition and prune the input domain to either uncover counterexamples or prove correctness. Prominent tools in this area, including $\alpha,\beta$-CROWN (Zhou et al., 2024), PyRAT (Lemesle et al., 2024), and Marabou (Wu et al., 2024), have demonstrated strong results in recent VNN-COMP competitions Brix et al. (2024), consistently ranking at the top of evaluation benchmarks. A key component of branch-and-bound approaches is computing tight output bounds for neural networks over constrained input regions. Techniques like interval arithmetic (Cheng et al., 2017; Pulina and Tacchella, 2010) are commonly employed to compute these bounds, influencing both the accuracy and computational efficiency of the verification.
@@ 2 Preliminaries
Notations. We write $x\sim\mathcal{X}$ to indicate that the random variable $x$ is drawn from the distribution $\mathcal{X}$, and $x\stackrel{{\scriptstyle\mathcal{X}}}{{\sim}}\mathbf{X}$ to indicate that $x$ is sampled from the set $\mathbf{X}$ according to the distribution $\mathcal{X}$. We use bold letters to represent sets and calligraphic letters to denote distributions. Consider a 3-dimensional tensor $x\in\mathbb{R}^{h\times w\times nc}$, which we flatten into a vector. The vectorized form is denoted as $\mathbf{vec}(x)\in\mathbb{R}^{n_{0}}$, where $n_{0}=nc\times w\times h$. The Minkowski sum is denoted by $\oplus$. The ceiling operator $\lceil x\rceil$ denotes the smallest integer greater than or equal to $x\in\mathbb{R}$. A feedforward neural network with input dimension $n_{0}$, $l$ hidden layers of sizes $n_{1},\ldots,n_{l}$, and output dimension $N$, using the $\mathrm{ReLU}$ activation function, is denoted as $n_{0},n_{1},\ldots,n_{l},N](\mathrm{ReLU})$. Her we denote, a $\ell_{2}$ ball with radius $r\in\mathbb{R}^{n_{0}}$ and center $x\in\mathbb{R}^{n_{0}}$ with $\mathbf{B}_{r}(x)$.
Semantic Segmentation Neural Networks
A semantic segmentation neural network (SSN) is a nonlinear function that assigns a class label to each pixel $x(i,j)$ in a multichannel input image $x$, producing a target class label $\mathsf{mask}(i,j)$ from a predefined set of classes $\mathbf{L}=\left\{1,2,\ldots,L\right\}$,
$\mathsf{SSN}:x\in\mathbb{R}^{h\times w\times nc}\mapsto\mathsf{mask}\in\mathbf{L}^{h\times w},$
where $h$, $w$, and $nc$ denote the height, width, and number of channels of the input image, respectively, and $(i,j)\in{1,\ldots,h}\times{1,\ldots,w}$ are pixel coordinates. In this paper, we denote the logits of SSNs by $y\in\mathbb{R}^{h\times w\times L}$, where the target class label $\mathsf{mask}(i,j)$ is given by $\mathsf{mask}(i,j)=\underset{l\in\mathbf{L}}{\arg\max}\ \ y(i,j,l)$. Here, we also define the function $f:\mathbb{R}^{n_{0}}\mapsto\mathbb{R}^{n}$ as a nonlinear map between the flattened images and the flattened logits. Formally,
[…]
Conformal Inference
Consider a collection of i.i.d and positive scalar random variables $\mathbf{M}=\left\{R_{1},R_{2},\ldots,R_{m}\right\}$ sampled from $R\sim\mathcal{D}$, where $R_{1}<R_{2}<\ldots<R_{m}$. Given a new draw $R_{m+1}$ from the same distribution and a miscoverage level $\epsilon\in(0,1)$, conformal inference (CI)(Vovk et al., 2005) constructs a prediction interval $C(R_{m+1})=[0,d]$ such that $\Pr[\ \ R_{m+1}\in C(R_{m+1})\ \ ]\ \geq 1-\epsilon$. In this paper, we refer to these scalar random variables as nonconformity scores and we refer to $\mathbf{M}$ as the calibration dataset. To compute the threshold $d$, we use the empirical distribution of nonconformity scores $R_{i}$ for $i\in 1,2,\ldots,m$, and define a rank $\ell:=\lceil(m+1)(1-\epsilon)\rceil$. A valid choice for the threshold $d$ is then $R_{\ell}$, the $\ell$-th smallest element of the augmented set $R_{1},\ldots,R_{m},\infty$}, as defined in (Tibshirani et al., 2019), which yields the following marginal guarantee:
@@
where the probability is over the joint randomness of both the calibration and test samples (Tibshirani et al., 2019; Vovk et al., 2005). The test sample $R_{m+1}$ typically refers to an unseen example drawn from the same distribution for which we aim to provide a coverage guarantee. In this paper, we also refer to this sample as the unseen, and denote it by $R^{\text{unseen}}$.
Given a sampled calibration dataset, the coverage level $\delta=\Pr[R^{\text{unseen}}\leq R_{\ell}]$ is itself a random variable following a $\mathrm{Beta}(\ell,m+1-\ell)$ distribution (Angelopoulos and Bates, 2021). The following equations shows the mean value and the variance of the beta distribution in terms of $m,\ell$:
$\mathbf{E}\big{[}\delta\big{]}=\frac{\ell}{m+1},\ \ \mathbf{Var}\big{[}\delta\big{]}=\frac{\ell(m+1-\ell)}{(m+1)^{2}(m+2)},$ (2)
that shows by appropriately tuning $m$ and $\ell$, we can significantly reduce the variance $\mathbf{Var}[\delta]$, to avhieve tighter bounds. For instance, when $m=8000$ and $\ell=7999$, the variance of the corresponding Beta distribution is $3.123\times 10^{-8}$, which is extremely small.
However, to present this in a completely formal way, we only considers low variance scenarios ($\mathbf{Var}(\delta)\ll 1$) and we also include this source of uncertainty by reformulating this guarantee with a double-step probablistic guarantee. To include this uncertainty, we can utilize the cumulative density function, CDF, of beta distribution- a function known as regularized incomplete beta function- and propose the following guarantee based on the definition of the CDF.
$\Pr\left[\ \ \Pr[\ \ R^{\text{unseen}}\leq R_{\ell}\ \ ]>1-\epsilon\ \ \right]>1-\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell).$ (3)
where $\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell)$ is the regularized incomplete beta function that is characterized by the choice of $m$ and $\ell$ evaluated at point $\delta_{1}=1-\epsilon$. The one-step marginal guarantee in (1) holds only when the rank is set to $\ell:=\lceil(m+1)(1-\epsilon)\rceil$. However, as shown in Appendix A, a key advantage of the double-step probabilistic guarantee is its flexibility: the strict relationship among $m$, $\ell$, and $\epsilon$ can be relaxed. Therefore, in this paper, we focus on the double-step guarantee and do not enforce the constraint $\ell:=\lceil(m+1)(1-\epsilon)\rceil$. Instead, we tune $m$, $\ell$, and $\epsilon$ such that $\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell)$ becomes negligibly small. Henceforth, for any logical statement $P(\ell,m)\in\left\{\mathrm{true},\mathrm{false}\right\}$, defined over the hyper-parameters $m,\ell\leq m$, we refer to the double-step guarantee,
@@
as the $\langle\epsilon,\ell,m\rangle$ guarantee and we say the statement $P(\ell,m)$ satisfies a $\langle\epsilon,\ell,m\rangle$ guarantee. We also refer to $\epsilon$ as the miscoverage level, $\delta_{1}:=1-\epsilon$, as the coverage level and $\delta_{2}:=1-\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell)$ as the confidence of guarantee. Appendix B provides a clarifying example on this topic.
Determinitic and Probabilistic Reachability Analysis on Neural Networks
Given a neural network $f:\mathbb{R}^{n_{0}}\rightarrow\mathbb{R}^{n}$ and an input set $\mathbf{I}\subset\mathbb{R}^{n_{0}}$, deterministic reachability analysis aims to construct a set $\mathbf{R}_{f}(\mathbf{I})\subset\mathbb{R}^{n}$ such that
@@
In contrast, probabilistic reachability analysis assumes a distribution $\mathcal{W}$ over the input set $\mathbf{I}$, denoted $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$, and for a given miscoverage level $\epsilon$, proposes a set $\mathbf{R}_{f}^{\epsilon}(\mathcal{W})\subset\mathbb{R}^{n}$ such that
@@
In this paper, we detail how to compute a probabilistic reach set with miscoverage level $\epsilon$, and sampling distribution $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$ by constructing a suitable calibration set $\mathbf{M}$ of size $m$, and selecting an appropriate rank $\ell$. Therefore to keep the terminlogies consistent, we reformulate the probabilistic reach set in (5) using the $\langle\epsilon,\ell,m\rangle$ guarantee, and we denote it by $\mathbf{R}_{f}^{\epsilon}(\mathcal{W}\ ;\ell,m)$, which satisfies
@@
An adversarial attack on an image involves perturbing the input using a set of noise images $x^{\text{noise}}_{1},\ldots,x^{\text{noise}}_{r}$ and corresponding coefficients $\lambda=[\lambda(1),\ldots,\lambda(r)]^{\top}$. These perturbations are applied through a parameterized function $\Delta_{\lambda,x^{\text{noise}}}(\cdot)$, generating the adversarial image as:
$x^{\text{adv}}=\Delta_{\lambda,x^{\text{noise}}}(x)=x+\sum_{i=1}^{r}\lambda(i)x_{i}^{\text{noise}}.$ (6)
We focus on evaluating the robustness of semantic segmentation networks (SSNs) under such adversarial settings, particularly under the unknown, bounded adversarial examples ($x^{\text{adv}}\in\mathbf{I}$) where the coefficients in vector $\lambda$ are unknown but bounded within specified ranges, i.e., $\lambda\in[\underline{\lambda},\bar{\lambda}]\subset\mathbb{R}^{r}$.
A pixel $x(i,j)$ is deemed attacked if it faces perturbation. It is also considered robust if, for all adversarial examples, the predicted label remains unchanged:  $\mathsf{SSN}(\Delta_{\lambda,x^{\text{noise}}}(x))(i,j)=\mathsf{SSN}(x)(i,j).$
Consider another scenario where there exists perturbations such that the label at $(i,j)$ changes. In this case, the pixel is non-robust if for all perturbations we have $\mathsf{SSN}(\Delta_{\lambda,x^{\text{noise}}}(x))(i,j)\neq\mathsf{SSN}(x)(i,j)$. The pixel is also unknown if for some perturbations we have $\mathsf{SSN}(\Delta_{\lambda,x^{\text{noise}}}(x))(i,j)=\mathsf{SSN}(x)(i,j)$, and for some other perturbations we have $\mathsf{SSN}(\Delta_{\lambda,x^{\text{noise}}}(x))(i,j)\neq\mathsf{SSN}(x)(i,j)$.
To quantify robustness of SSN we consider two different metrics:
Robustness Value.: The Robustness Value (RV) of a network is the percentage of pixels that remain robust under attack: $\quad RV=100\times(N_{\text{robust}}/N_{\text{pixels}}),\quad\text{where }N_{\text{pixels}}=h\times w$.

### v2

@@ Abstract
Semantic segmentation networks (SSNs) are central to safety-critical applications such as medical imaging and autonomous driving, where robustness under uncertainty is essential. However, existing probabilistic verification methods often fail to scale with the complexity and dimensionality of modern segmentation tasks, producing guarantees that are overly conservative and of limited practical value. We propose a probabilistic verification framework that is architecture-agnostic and scalable to high-dimensional input-output spaces. Our approach employs conformal inference (CI), enhanced by a novel technique that we call the clipping block, to provide provable guarantees while mitigating the excessive conservatism of prior methods. Experiments on large-scale segmentation models across CamVid, OCTA-500, Lung Segmentation, and Cityscapes demonstrate that our framework delivers reliable safety guarantees while substantially reducing conservatism compared to state-of-the-art approaches on segmentation tasks. We also provide a public GitHub repository [^1] https://github.com/Navidhashemicodes/SSN_Reach_CLP_Surrogate for this approach, to support reproducibility.
@@ 1 Introduction
Semantic image segmentation is widely applied in high-stakes areas, including medical imaging (perone2018spinal) and autonomous driving (deng2017cnn). Yet, segmentation models powered by deep learning are known to be fragile, as they can be easily manipulated by adversarial inputs (xie2017adversarial), limiting their safe use in these critical applications.
[…]
A prominent research direction in probabilistic verification is randomized smoothing, first introduced for classifiers cohen2019certified and later adapted to segmentation models fischer2021scalable, with subsequent improvements in anani2024adaptive; hao2022gsmooth. More recently, conformal inference has been investigated as a means to provide probabilistic guarantees in classification tasks jeary2024verifiably. In hashemiscaling, the authors proposed a methodology to extend the application of conformal inference from robustness analysis in classification task to robustness analysis in segmentation tasks. They first introduce a naive approach that relies solely on conformal inference, which produces only hypercubes for the reachable set of the vector of logits—also described in the preliminaries of this paper. They then improve this method by training a small ReLU neural network as a surrogate for the base segmentation model, demonstrating that the surrogate enables the construction of more general convex shapes for the system’s reachable set.
Appendix A presents a comprehensive comparison between segmentation verification techniques based on randomized smoothing fischer2021scalable; anani2024adaptive and the conformal inference approach in hashemiscaling, highlighting the advantages of conformal inference for segmentation tasks. However, as we detail later, the method in hashemiscaling has its own limitations, which we address in this paper. In fact, the small ReLU neural network introduced in hashemiscaling poses several challenges for this methodology. First, the surrogate model can suffer from fidelity issues, leading to overly conservative guarantees. Second, the high dimensionality of the input space makes training infeasible. Moreover, it imposes restrictive assumptions on the perturbation of the image, requiring it to be sparse and confined to $\ell_{\infty}$-type perturbations.
In this paper, we propose an alternative technique to replace this ReLU surrogate model and to avoids all these issues. Specifically, we introduce a clipping-block at the end of the base model. This approach requires no surrogate training—eliminating fidelity concerns—and allows perturbations across the entire image, supporting general $\ell_{p}$ perturbations. To show this numerically, we deliberately consider the same models and datasets as in hashemiscaling, but with higher-dimensional (and also full) image perturbations in our numerical experiments and we observed that the surrogate-based approach in hashemiscaling fails under these conditions because deterministic reachability using the NNV toolbox tran2020nnv cannot handle the dimensionality of the input sets used in our experiments facing runtime and out-of-memory errors.
@@
Deterministic Verification of Neural Networks. Deterministic verification of neural networks has been addressed using several well-established techniques, such as Satisfiability Modulo Theories (SMT) solvers (duong2023dpll; katz2017reluplex; wu2024marabou), Mixed Integer Linear Programming (MILP) formulations (anderson2020strong; cheng2017maximum; tjeng2017evaluating), and different types of reachability analysis (bak2020improved; tran2020verification). Many of the existing methods can be also viewed through the general framework of branch-and-bound algorithms (bunel2020branch), which iteratively partition and prune the input domain to either uncover counterexamples or prove correctness. Prominent tools in this area are, $\alpha,\beta$-CROWN, (zhou2024scalable), PyRAT (lemesle2024neural), and Marabou (wu2024marabou). Existing deterministic techniques do not scale to segmentation models because of their high level of complexity and dimensionality.
Conformal Inference. The integration of CI with formal verification techniques has recently received noticeable interest, that is primarily due to its accuracy and level of scalability. For instance, bortolussi2019neural merges CI with neural state classifiers to develop a stochastic runtime verification algorithm. lindemann2023safe; zecchin2024forking employs CI to guarantee results in MPC control using a trained model. tonkens2023scalable applies CI for planning with probabilistic safety guarantees, and hashemi2023data; hashemi2024statistical employs CI for reachability of stochastic dynamical systems, see lindemann2024formal for a recent survey article.
@@ 2 Preliminaries
Notations. We write $x\sim\mathcal{X}$ to indicate that the random variable $x$ is drawn from the distribution $\mathcal{X}$, and $x\stackrel{{\scriptstyle\mathcal{X}}}{{\sim}}\mathbf{X}$ to indicate that $x$ is sampled from the set $\mathbf{X}$ according to the distribution $\mathcal{X}$. We use $\mathsf{betacdf}$ to denote the CDF of beta distribution. We use bold letters to represent sets and calligraphic letters to denote distributions. Consider a 3-dimensional tensor $x\in\mathbb{R}^{h\times w\times nc}$ of height $h$, width $w$, and $nc$ channels, which we flatten into a vector. The vectorized form is denoted as $\mathbf{vec}(x)\in\mathbb{R}^{n_{0}}$, where $n_{0}=nc\times w\times h$. The Minkowski sum is denoted by $\oplus$. The ceiling operator $\lceil x\rceil$ denotes the smallest integer greater than or equal to $x\in\mathbb{R}$. Here, we denote a $\ell_{2}$ ball with radius $r\in\mathbb{R}^{n_{0}}$ and center $x\in\mathbb{R}^{n_{0}}$ with $\mathbf{B}_{r}(x)$.
Semantic Segmentation Neural Networks (SSN)
Consider an input image $x\in\mathbb{R}^{h\times w\times nc}$. An SSN processes this image and outputs a $2$-D label mask, assigning a class to every pixel location $(i,j)\in\left\{1,\ldots,h\right\}\times\left\{1,\ldots,w\right\}$, we have: $\text{SSN}(x)=\text{mask}\in\mathbf{L}^{h\times w},\quad\mathbf{L}=\{1,2,\ldots,L\}$. The network computes the logits $y\in\mathbb{R}^{h\times w\times L}$ in the last layer. For each pixel, the predicted label is obtained by choosing the class with the largest logit value: $\text{mask}(i,j)=\arg\max_{\ell\in L}\;y(i,j,\ell).$
Instead of working directly with multidimensional tensors, it is sometimes convenient to flatten both the input image and the logits. Let $n_{0}=h\cdot w\cdot nc$ and $n=h\cdot w\cdot L$. Then the map between flattened input $\mathbf{vec}(x)$ and the flattened logit values $\mathbf{vec}(y)$ can equivalently be viewed as a nonlinear function
$f:\mathbb{R}^{n_{0}}\to\mathbb{R}^{n},\quad f(\text{vec}(x))=\text{vec}(y),$
Conformal Inference and $\langle\epsilon,\ell,m\rangle$ guarantee
Suppose we have $m$ i.i.d. non-negative random variables $\mathbf{M}=\{R_{1},R_{2},\ldots,R_{m}\},\quad R_{1}<R_{2}<\cdots<R_{m}$, drawn from some distribution $R\sim\mathcal{D}$. Conformal inference (CI) (vovk2005algorithmic) provides a way to construct prediction intervals with guaranteed coverage at a user-specified miscoverage level $\epsilon\in(0,1)$.
For a new draw $R_{m+1}$ from the same distribution $R_{m+1}\sim\mathcal{D}$, CI defines a prediction interval $C(R_{m+1})=[0,d]$ such that $\Pr\big[R_{m+1}\in C(R_{m+1})\big]\geq 1-\epsilon$.
In this setting, the values $R_{i}\in\mathbf{M}$ are often called nonconformity scores, and the set $\mathbf{M}$ is referred to as the calibration dataset. To compute the threshold $d$, one considers the empirical distribution of these scores. Specifically, define $\ell=\lceil(m+1)(1-\epsilon)\rceil$, and let $R_{\ell}$ denote the $\ell$-th smallest element ($\ell<m$) of the set $\mathbf{M}$ (tibshirani2019conformal). Then the prediction threshold is given by $d=R_{\ell}$, and we are guaranteed the following marginal coverage guarantee:
@@
Here, the probability is marginal which means that it is taken jointly over both the randomness of calibration dataset $\mathbf{M}$ and the test sample $R_{m+1}$ (tibshirani2019conformal; vovk2005algorithmic). The test sample $R_{m+1}$ is assumed to be drawn from the same underlying distribution as the calibration set. Following the convention in this paper, we call this new sample unseen and denote it by $R^{\text{unseen}}$. Since we do not assume randomness in the calibration dataset, following the proposition of vovk2005algorithmic the authors in hashemiscaling use a two-step representation of the guarantee:
$\Pr\left[\ \ \Pr[\ \ R^{\text{unseen}}\leq R_{\ell}\ \ ]>1-\epsilon\ \ \right]>1-\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell),$ (2)
where the outer probability captures the randomness in $\mathbf{M}$, while the inner probability is valid for a fixed set $\mathbf{M}$. For further details and clarifying examples, please see hashemiscaling.
Here, we tune $m$, $\ell$, such that for a given $\epsilon\in[0,1]$, $\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell)$ becomes less than a user-specified threshold. Henceforth, for any logical statement $P(\ell,m)\in\left\{\mathrm{true},\mathrm{false}\right\}$, defined over the hyper-parameters $m,\ell\leq m$, we refer to the double-step guarantee,
@@
as the $\langle\epsilon,\ell,m\rangle$ guarantee and we say $P(\ell,m)$ satisfies a $\langle\epsilon,\ell,m\rangle$ guarantee. We refer to $\delta_{1}:=1-\epsilon$, as the coverage level and $\delta_{2}:=1-\mathsf{betacdf}_{1-\epsilon}(\ell,m+1-\ell)$ as the confidence level.
Deterministic and Probabilistic Reachability Analysis on Neural Networks
Reachability analysis for a neural network $f:\mathbb{R}^{n_{0}}\rightarrow\mathbb{R}^{n}$ has been approached in the literature in two ways. In deterministic reachability analysis, given an input set $\mathbf{I}\subset\mathbb{R}^{n_{0}}$, the goal is to construct a set $\mathbf{R}_{f}(\mathbf{I})\subset\mathbb{R}^{n}$ that contains all possible outputs:
@@
Probabilistic reachability analysis, in contrast, assumes a prior distribution $\mathcal{W}$ for inputs over $\mathbf{I}$, denoted $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$. For a given miscoverage level $\epsilon$, it aims to construct a set $\mathbf{R}_{f}^{\epsilon}(\mathcal{W})\subset\mathbb{R}^{n}$ such that the network output lies within this set with high probability:
@@
In this work, we detail a procedure to compute such a probabilistic reach set for a given miscoverage level $\epsilon$ and sampling distribution $x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$. This involves constructing a calibration set $\mathbf{M}$ of size $m$ and selecting a rank $\ell$ to enforce the desired coverage. To maintain consistent terminology, we express the probabilistic reach set using the $\langle\epsilon,\ell,m\rangle$ guarantee and denote it as $\mathbf{R}_{f}^{\epsilon}(\mathcal{W}\ ;\ell,m)$, which satisfies
@@
Adversarial attack is a perturbation on an image by adding a combination of noise images $x^{\text{noise}}_{1},\ldots,x^{\text{noise}}_{r}$ weighted by a vector of coefficients $\lambda=[\lambda(1),\ldots,\lambda(r)]^{\top}$. Formally, an adversarial image is generated via
$x^{\text{adv}}=\Delta_{\lambda,x^{\text{noise}}}(x)=x+\sum_{i=1}^{r}\lambda(i)x_{i}^{\text{noise}},\quad\lambda\in[\underline{\lambda},\bar{\lambda}]\subset\mathbb{R}^{r},$ (5)
where the coefficients are unknown but bounded, defining the set of feasible adversarial inputs $x^{\text{adv}}\in\mathbf{I}$. The impact of these perturbations on semantic segmentation networks (SSNs) can be analyzed at the pixel level.
The SSN at pixel (i,j) is robust if its predicted label is invariant under all allowed perturbations: $\mathsf{SSN}(\Delta_{\lambda,x^{\text{noise}}}(x))(i,j)=\mathsf{SSN}(x)(i,j),\quad\forall\lambda\in[\underline{\lambda},\bar{\lambda}]$. If the label changes under all perturbations, the pixel is non-robust: $\mathsf{SSN}(\Delta_{\lambda,x^{\text{noise}}}(x))(i,j)\neq\mathsf{SSN}(x)(i,j),\quad\forall\lambda\in[\underline{\lambda},\bar{\lambda}]$. A pixel is classified as unknown when some perturbations change its label while others do not.
[…]
Once the reachset $\mathbf{R}_{f}^{\epsilon}(\mathcal{W}\ ;\ell,m)$ over the SSN logits $y\in\mathbb{R}^{n}$ is obtained, we can analyze each pixel status by projecting it onto individual logit components. This projection yields, for every pixel location $(i,j)\in\{1,\ldots,h\}\times\{1,\ldots,w\}$, a set of $L$ intervals corresponding to the class labels $l\in\mathbf{L}$.
For pixel $(i,j)$, let $l^{*}(i,j)\in\mathbf{L}$ denote the class whose logit interval has the largest lower bound. The pixel is classified as unknown if this lower bound does not strictly exceed the upper bounds of all other class intervals, indicating that multiple labels are possible under perturbation. Otherwise, if the lower bound of $l^{*}(i,j)$ is strictly greater than all others, the pixel is labeled based on the original prediction: it is robust if $l^{*}(i,j)=\mathsf{SSN}(x)(i,j)$ and non-robust if $l^{*}(i,j)\neq\mathsf{SSN}(x)(i,j)$. Algorithm 2 provides a detailed step-by-step procedure for computing the pixel-wise status using this approach.